OpenStudy (anonymous):

What is the equation of the line, in point-slope form, that passes through the points (-3, -1) and (-6, 8)?

OpenStudy (anonymous):

A. y + 8 = -3(x - 6) B. y - 8 = -3(x + 6) C. y + 8 = 3(x - 6) D. y - 8 = 3(x + 6)